# SpIDER: Spatially Informed Dense Embedding Retrieval for Software Issue Localization

### v1

@@ 4 Building Code Graphs
While the graph includes four types of nodes, we focus on retrieving functions, which represent the most granular and therefore the most challenging module type. This choice is further motivated by data from the SWE-PolyBench dataset, where the majority of instances ($67\%$) contain ground truth edits at the function level, with an average of $2.78$ edits per instance rashid2025swepolybenchmultilanguagebenchmarkrepository. In contrast, only $1.42\%$ of instances involve edits at the class-only level (i.e., excluding changes to methods within the class), with an average of just $0.49$ edits per instance. This disparity highlights both the prevalence and complexity of function-level edits, highlighting the importance of accurate function-level retrieval. Additionally, since files are composed of functions and classes, effective retrieval at the function and class level naturally supports accurate file-level retrieval as well. Table 1 shows the final dataset statistics. We note down the percentage of instances retained from each dataset for all languages. The retention depends on the type of ground truth node (class, file or function) that needs to be edited to fix the issue for a given instance. If any of the ground truth nodes for an instance is a function, then the instance is retained and ignored otherwise. We also report the average edits per instance for all the languages and datasets.
5 SpIDER Framework
Algorithm 1 SpIDER
0: Code graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, issue description text $Q$, code/text dense encoder $\mathcal{F}(\cdot)$,Parameters: retrieval budget $K$, filtering threshold NN, number of centers $C$, search depth $d$
1: Compute node embeddings: $\mathcal{F}(\mathcal{V})=\{\mathcal{F}(v_{i})\}_{i=1}^{n}$
2: Compute query embedding: $\mathcal{F}(Q)$
3: Compute cosine similarities: $\mathcal{S}=\{s_{i}=f(v_{i})\cdot f(Q)\}_{i=1}^{n}$
4: Let $\mathcal{I}_{K}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$K$ nodes sorted by $s_{i}$ in descending order
5: Let $\mathcal{I}_{N}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$N$ nodes sorted by $s_{i}$ in descending order, with $N>K$
6: Let $\mathcal{V}_{K}=\{v_{i}\mid i\in\mathcal{I}_{K}\}$ and $\mathcal{V}_{N}=\{v_{i}\mid i\in\mathcal{I}_{N}\}$
7: Let $\mathcal{I}_{C}\subset\mathcal{I}_{K}$ be the set of first $C$ indices in $\mathcal{I}_{K}$
8: Let $\mathcal{V}_{C}=\{v_{i}\mid i\in\mathcal{I}_{C}\}$ {Center nodes for BFS}
9: For each node $v\in\mathcal{V}_{C}$, perform BFS in $\mathcal{G}$ up to depth $d$ and collect neighbors of type ‘function‘ that also belong to $\mathcal{V}_{N}$
10: Apply an LLM-based filtering to these neighbors and retain those most likely to resolve the issue $Q$. The LLM is provided the source code content of all the neighbors of a center node along with the issue description and asked to select only the neighbors that are most likely to solve the issue. Let $\mathcal{I}_{\text{LLM}}$ denote the set of indices selected by the LLM, consisting of the indices of the center nodes and the LLM-selected neighbors from among the BFS retrieved neighbors.
[…]
Figure 1 and Algorithm 1 provides a high-level overview of our proposed framework. Given an issue description (bug report, feature requests, etc.) and codebase, SpIDER retrieves top $K$ most relevant functions that should be modified to solve the issue. This method can be easily extended to class level as well as file level retrieval by using subset of code chunks if the entire class or file level code cannot fit into the LLM context (refer to section˜A.2 for more details). For this reason, we use the terms ‘function retrieval’ and ‘node retrieval’ interchangeably.
SpIDER’s design reflects our key observation: for instances with multiple edits, it is often the case that the edits are spatially close to each other in a code graph due to the dependencies on other functions/classes/files. Therefore, modifying a function can have cascading effects, requiring modifications to all other functions that call it, as well as to its containing class/file. Similarly, while semantically retrieved nodes offer diverse coverage of the potential candidate functions across the codebase, they may not be fully leveraging the "near miss" scenarios, where the true solution may spatially close to the top $K$ ranked functions. SpIDER tries to improve the coverage by balancing the neighborhood exploration and semantic retrieval to retrieve the most likely set of $K$ functions.
Semantic Retrieval.
The first step is to retrieve the top $K$ functions based on the semantic similarity of their dense embeddings to the issue description. We can obtain dense embeddings from encoders like CodeSAGE zhang2024codesage or an embedding model from SweRank reddy2025sweranksoftwareissuelocalization. Once function embeddings are generated in the code graph, we rank functions by cosine similarity to the issue description embedding. Next, we retrieve top $K$ functions from the semantically ranked list of functions. We call the set of node indices of the top $K$ functions ranked in the descending order of their similarity scores with the issue descriptions as $\mathcal{I}_{K}$. We also store a set of node indices of top $N$ ranked functions $\mathcal{I}_{N}$ where $N>K$. This is our primary set of retrieved functions which we aim to further improve using neighborhood exploration.
Neighborhood Exploration.
Since we have a retrieval limit of $K$, it is not affordable to store the entire neighborhood of all the nodes in $\mathcal{I}_{K}$. To strictly limit the neighbors we explore and add to the final retrieval list of the same size $K$, we restrict the neighborhood exploration to the first $C$ nodes in $\mathcal{I}_{K}$ within a $d$-hop distance from the center nodes. While we primarily focus on exploring neighbors via the ‘contains’ edge type to capture the structural organization of the codebase, SpIDER can be extended to incorporate other edge types like ‘imports’, ‘invokes’, ‘inherits’, etc. for further performance enhancements. Two functions within the same class would be 2 hops away from each other. Similarly, two functions within two separate classes in two separate files will be 4 hops away. We do not explore neighbors beyond $d$ hop distance. Increasing the depth improves coverage at the cost of increased neighbors context per center. Hence, $d$ and $C$ are hyperparameters that a user can adjust based on the downstream budget, thereby offering more flexibility and tuning capabilities in terms of retrieval performance.
For evaluation, we set $C=5$ and $N=500$ based on optimal performance observed for Python in the SWE-PolyBench dataset (see table˜4 and table˜12). These values are kept fixed across all languages and datasets to avoid hyperparameter overfitting. Additionally, we set $d=4$ to ensure that token consumption per LLM invocation remains below 10 00010\,000. While higher values of $K$ generally improve retrieval performance (see table˜13, fig.˜2), we evaluate for $K=20$ in our main results (Tables 3 and 2) to ensure fairness to the SweRank baseline, which selects $K=20$ as optimal for its downstream LLM-based reranker that improves localization at higher ranks.
Neighbor filtering
[…]
6 Experiments
Our goal is to answer the following three research hypotheses from our experiments:
H1 Can the graph structure of a codebase be leveraged to improve retrieval recall and accuracy for instances involving multiple function edits that are spatially close within the code graph?
H2 Among the top-$K$ nodes retrieved using dense embedding retrieval, do the highest-ranked nodes serve as accurate approximations of the actual ground truth locations in the code graph? If so, is it possible to reach the true nodes through local exploration of the code graph, potentially trading off some lower-ranked retrieved nodes in favor of better spatial coverage?
H3 How do existing retrieval methods, along with our proposed approach, perform across programming languages beyond Python?
To this end, we focus our study towards scenarios having more than 1 edit per instance as shown in table 1 for 4 programming languages, namely Python, Java, JavaScript and TypeScript, highlighting the robustness of the baselines and our approach.
@@ 6.1 Experimental Setup
(a)  Python
(b)  Java
(c)  JavaScript
(d)  TypeScript
Figure 2: Dense retrieval methods performance comparison on SWE-PolyBench benchmark.
Datasets
@@
We conduct experiments on SWEBench-Verified, SWE-PolyBench and Multi-SWEBench considering patched functions as the ground truth targets for retrieval. Following suresh2025cornstack, we exclude the instances having no function level modifications in their patch. We evaluate across the Python, Java, TypeScript, and JavaScript programming languages.
Evaluation metrics
Since, the goal is to improve the coverage when retrieving $K$ nodes, we mainly use Recall@$K$ and Acc@$K$ for evaluating and comparing various methods. Here Acc@$K$ for an instance is 1 if all the ground truth targets are present in $K$ retrieved predictions. We also present Mean Reciprocal Rank (MRR) which is the inverse of the rank of the first encounter of the ground truth node in the top-$K$ retrieved list.
Hyperparameters
The hyperparameters for our algorithm are $C$, $d$, NN and $K$. We present various studies ablating on each of these parameters in the subsequent sections. Following reddy2025sweranksoftwareissuelocalization; fehr-etal-2025-coret we use $K=20$ for all the datasets to be fair to localization methods that are sensitive to the choice of this hyperparameter such as SWERankEmbed + reranker. We perform a grid search to select SpIDER’s hyperparameters on SWE-PolyBench’s Python split, resulting in $C=5$, $d=4$ and $N=500$. We keep these values constant for all our experiments. In our ablation studies, we analyze how these design choices affect the LLM’s context length requirements and invocation budget, providing deeper insights for downstream users. For all experiments, we set the LLM temperature to $0.1$.
Baselines
We compare SpIDER against state-of-the-art retrieval approaches based on dense embedding models like SweRankEmbed-Small ZS, CodeSAGE-Small ZS, reddy2025sweranksoftwareissuelocalization; zhang2024codesage and BM25 Robertson1994OkapiATBM25 as it represent a popular sparse retrieval method. Because CodeSAGE-Small ZS is not trained for Github-style problem solving tasks, we further finetune it using parameter efficient finetuning using LoRA finetuning (hu2022lora) on $2000$ instances from 3 repositories of the SWEBench training dataset. We call this baseline CodeSAGE-Small-PEFT. fehr-etal-2025-coret did not release the embedding model as well as code publicly. Hence, we could not use it as a baseline. We do not include LocAgent (chen2025locagent) due to two main drawbacks: it is extremely slow and requires significantly more number of LLM invocations as opposed to SpIDER. To give an estimate of performance of LocAgent compared to other methods, we evaluate its performance on SWEBench-Verified and on Python repositories of SWE-PolyBench benchmark table˜9. It took us more than 50 hours to complete the evaluation of LocAgent on SWEBench-Verified while SpIDER finished the evaluation in less than 2 hours. Additionally, SpIDER requires $C$ invocations of an LLM which is not deterministic for LocAgent.
We improve each baseline using spatial information from code graphs and demonstrate two retrieval approaches for each model, one using SpIDER and the other using dense retrieval with cosine similarity for dense embedding models and the standard sparse retrieval implementation for BM25.
@@ 6.2 Results
Table 2: Comparison of dense retrieval and dense retrieval + reranker/selector for $K=20$ and $N=500$, $C=5$, $d=4$. Winners per metric are highlighted in light blue.
Language Model Retrieval Dense Retrieval Only Dense + Reranker/Selector
Recall@20 Accuracy@20 MRR@20 Recall@3 Accuracy@3 MRR@3
Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.30 0.44 0.36 0.54
DER 0.42 0.35 0.26 0.38 0.31 0.48
CodeSAGE-Small ZS SpIDER 0.27 0.21 0.13 0.23 0.18 0.31
DER 0.21 0.17 0.11 0.19 0.16 0.25
CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.24 0.32 0.25 0.42
DER 0.29 0.20 0.21 0.26 0.20 0.38
BM25 SpISR 0.33 0.26 0.19 0.30 0.23 0.38
SR 0.31 0.24 0.18 0.27 0.22 0.35
Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.18 0.25 0.17 0.36
DER 0.31 0.24 0.18 0.24 0.16 0.33
CodeSAGE-Small ZS SpIDER 0.23 0.17 0.10 0.16 0.09 0.26
DER 0.15 0.10 0.07 0.11 0.07 0.19
CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.15 0.20 0.14 0.28
DER 0.21 0.14 0.13 0.17 0.11 0.25
BM25 SpISR 0.16 0.10 0.10 0.12 0.08 0.20
SR 0.15 0.10 0.10 0.12 0.09 0.19
DER 0.52 0.45 0.16 0.41 0.34 0.42
CodeSAGE-Small ZS SpIDER 0.38 0.32 0.09 0.30 0.25 0.31
DER 0.26 0.21 0.07 0.22 0.17 0.24
CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.15 0.41 0.35 0.41
DER 0.45 0.38 0.14 0.36 0.30 0.39
BM25 SpISR 0.46 0.39 0.13 0.36 0.30 0.38
SR 0.36 0.29 0.11 0.31 0.25 0.34
DER 0.29 0.21 0.18 0.25 0.19 0.30
CodeSAGE-Small ZS SpIDER 0.16 0.10 0.03 0.14 0.09 0.19
DER 0.10 0.06 0.06 0.09 0.06 0.14
CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.04 0.22 0.16 0.27
DER 0.19 0.13 0.11 0.18 0.12 0.24
BM25 SpISR 0.11 0.09 0.02 0.10 0.07 0.11
SR 0.07 0.05 0.04 0.07 0.05 0.09
Leveraging Code Graphs Improves Retrieval
We assess the effect of graph structure on code retrieval in SWE-PolyBench, comparing standard dense embeddings to our graph-aware SpIDER across Python, TypeScript, JavaScript, and Java (Fig. 2). The top-$K$ modules are retrieved by ranking code embeddings against issue embeddings via cosine similarity. SpIDER (SweRankEmbed-Small + SpIDER), which incorporates spatially informed graph structure, consistently outperforms all baselines across languages and $K$ values, highlighting the benefit of leveraging codebase graph information. Notably, the zero-shot SweRankEmbed-Small model– trained solely on $\approx$3,300 Python repositories– remains competitive across non-Python languages, surpassing both BM25 and codesage-small-v2 in zero-shot and PEFT settings, demonstrating strong cross-lingual transfer. Overall, while semantic embeddings provide a solid foundation, explicitly modeling spatial relationships via SpIDER yields substantial additional gains in retrieval accuracy.
Table 3: Dense retrieval performance across benchmarks for $K=20$, $N=500$, $C=5$, $d=4$. Winners per metric highlighted in light blue.
Language Model Retrieval Recall@20 Accuracy@20 MRR@20

### v2

@@ 4 Methodology
4.1 Building Code Graphs
For a given codebase, we construct a code graph using only files written in the primary programming language, defined as the language comprising the majority of the repository. Files in other languages are ignored to simplify graph construction. For example, in a Python repository, we consider only Python source files.
We follow the graph schema proposed by Chen et al. (2025), where each node represents a code entity—function, class, file, or directory—and edges encode structural relations of types imports, invokes, contains, and inherits. For Python, we use the built-in ast module [^1] https://docs.python.org/3/library/ast.html to parse source files and extract syntax trees. For Java, JavaScript, and TypeScript, we rely on Tree-sitter [^2] https://github.com/tree-sitter for syntax parsing.
Graph construction proceeds by traversing the repository hierarchy. Directory nodes are added to enable navigation across the project structure. File nodes are included only if they contain at least one function or class definition; files containing solely metadata or configuration are ignored. Function and class nodes, along with the edges connecting them, are extracted from the corresponding syntax trees. While directory nodes do not store code content, all other node types include their associated source code.
Constructing accurate graphs for Java, JavaScript and TypeScript presents additional challenges due to the flexible ways in which functions and classes can be declared/implemented. For example, class definitions and method implementations may be split across multiple locations or files. We explicitly handle such language-specific cases using Tree-sitter to ensure faithful graph construction. Refer App. A.8.
Although the resulting graph contains multiple node types, we focus exclusively on retrieving functions, which constitute the most granular, and consequently the most challenging retrieval unit. This choice is motivated by statistics from the SWE-PolyBench dataset, where $67\%$ of instances involve function-level edits, with an average of $2.78$ edits per instance (Rashid et al., 2025). In contrast, only $1.42\%$ of instances involve class-only edits (excluding method changes), with an average of $0.49$ edits per instance. Since files are composed of functions and classes, accurate retrieval at finer granularities naturally helps file-level localization as well.
Table 1 reports dataset statistics, including the proportion of retained instances per language. For function-level retrieval, we exclude instances involving only file- or class-level edits. Similarly, for class- and file-level retrieval, we ignore instances with edits exclusively at other granularities.
4.2 Leveraging Code Graph via SpIDER
fig.˜1 provides a high-level overview of SpIDER, with algorithmic details presented in algorithm˜1 (section˜A.7). Given an issue description $Q$ and a codebase represented as graph $\mathcal{G}$, SpIDER retrieves the top-$K$ functions most likely to require modification. While we focus on function-level retrieval, the framework naturally extends to class- and file-level retrieval by aggregating subsets of code chunks when full contexts exceed the LLM budget (see section˜A.9).
SpIDER’s design is motivated by the observation that, in instances requiring multiple edits, relevant functions tend to be structurally proximate in the graph due to call dependencies and containment relationships. While dense semantic retrieval provides broad coverage, it often fails to capture “near-miss” cases, where the correct function lies close structurally but not semantically to top-ranked candidates.
Semantic Retrieval  We first compute the baseline dense retrieval $\mathcal{S}_{K}(Q)$ as defined in section˜3. Additionally, we compute $\mathcal{S}_{N}(Q)$ containing the top-$N$ nodes ($N>K$) to constrain neighborhood exploration.
Seed Selection  From $\mathcal{S}_{K}(Q)$, we select the top-$C$ ranked nodes as seed centers for graph exploration:
$\mathcal{C}_{Q}:={v\in\mathcal{S}_{C}(Q)}$
where $C\leq K$. A seed anchors a local neighborhood search.
Neighborhood Exploration  For each seed, we perform breadth-first search up to depth $d$ along ‘contains’ edges, collecting structurally proximate function nodes. Formally, the $d$-hop neighborhood is:
$\Gamma_{d}(\mathcal{C}_{Q}):=\{u\in\mathcal{V}\mid\mathrm{dist}_{\mathcal{G}}(u,\mathcal{C}_{Q})\leq d\}$
where $\mathrm{dist}_{\mathcal{G}}(u,\mathcal{C}_{Q})$ denotes the shortest-path distance from $u$ to any node in $\mathcal{C}_{Q}$. Two functions within the same class are 2 hops apart; functions in different classes across separate files are 4 hops apart. While we use ‘contains’ edges to capture hierarchical structure, the framework can incorporate other edge types like ‘invokes’, ‘imports’, and ‘inherits’.
Two-Stage Neighbor Filtering
To control computational cost while maintaining quality, candidates are filtered in two stages. Stage 1 (Semantic Filtering): We restrict neighbors to those also appearing in the top-$N$ candidate pool:
$\hat{\mathcal{C}}_{Q}:=\Gamma_{d}(\mathcal{C}_{Q})\cap\mathcal{S}_{N}(Q)$
This ensures exploration remains within semantically relevant regions of the graph. Stage 2 (LLM-based Selection): An LLM selector $\mathcal{L}$ evaluates each candidate $u\in\hat{\mathcal{C}}_{Q}$, receiving the source code content and issue description as context. The LLM returns a binary relevance decision $\mathcal{L}(u)\in\{0,1\}$. Note that seed centers $\mathcal{C}_{Q}$ are always retained in $\mathcal{U}_{Q}$, where:
$\mathcal{U}_{Q}:={u\in\hat{\mathcal{C}}_{Q}\mid\mathcal{L}(u)=1}\cup\mathcal{C}_{Q}$
Output Construction  To maintain a fixed retrieval budget of $K$, newly selected neighbors replace the lowest-ranked non-center nodes. The final retrieval set is:
$\mathcal{S}^{\texttt{LLM}}_{K}(Q):=\mathcal{U}_{Q}\cup\underset{v\in\mathcal{\mathcal{S}_{K}(Q)}}{\mathrm{arg\,top}_{K-|\mathcal{U}_{Q}|}}\;s_{Q}(v)$ (1)
HyperparametersSpIDER introduces four hyperparameters: retrieval budget $K$, number of seed centers $C$, exploration depth $d$, and semantic filtering threshold $N$. We analyze their effects in section˜5.3.
Definition 4.1 (Expected Recall Change).
$\pi_{B}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\mathcal{S}_{K}\setminus\mathcal{S}_{K-|\mathcal{U}_{Q}|}),$
$\pi_{\Gamma}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\hat{\mathcal{C}}_{Q}\setminus\mathcal{S}_{K}).$
Proposition 4.2 (Sufficient Condition for Recall Improvement).
Following the definitions from sections 3 and 4.2 and under the assumptions 1, 2 and 3 listed in section˜A.2, if it holds that $\alpha\pi_{\Gamma}>\pi_{B}+\beta(1-\pi_{\Gamma})$, then we can say,
$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]>\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})].$
Proposition 4.2 shows that when (i) relevant functions form a localized region in the code graph, (ii) dense retrieval places at least one seed in that region, and (iii) the LLM selector has higher true-positive than false-positive rate, then graph-aware retrieval strictly improves expected recall at fixed budget $K$. Proof can be found in section˜A.1.
5 Experiments and Results
We design our experiments to investigate whether leveraging structural relationships within code improves retrieval performance, particularly in challenging function-level settings. We focus on three high-level research questions:
RQ1 (Exploiting Graph Structure) Can incorporating the structural relationships between code modules improve retrieval of relevant functions beyond what semantic similarity alone achieves?
RQ2 (Cross-Language Robustness) Do retrieval strategies that leverage code structure generalize across programming languages and codebases of varying complexity?
RQ3 (Efficiency and Practicality) Can graph-aware retrieval methods achieve improved coverage without incurring excessive computational cost or LLM invocations?
@@ 5.1 Experimental Setup
Figure 2: Function-level retrieval performance on SWE-PolyBench. Across all programming languages and retrieval budgets $K$, SWERankEmbed-Small+SpIDER consistently outperforms dense and sparse baselines in both Accuracy and Recall, highlighting robust gains from incorporating graph-aware neighborhood expansion under a fixed retrieval budget. Shaded envelopes indicate standard deviations estimated via bootstrapping with $1000$ draws.
@@
Datasets  To answer our research questions, we focus on challenging multi-function edits and evaluate retrieval performance across multiple programming languages. Specifically, we conduct experiments on SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench, using patched functions as the ground-truth targets for retrieval. Following Suresh et al. (2025), we exclude instances that do not contain function-level modifications. The evaluation spans Python, Java, TypeScript, and JavaScript repositories, allowing us to assess both the robustness of our approach and the generalization of baseline methods across different languages and codebase complexities.
Evaluation Metrics  Given a query $Q$ and a retrieved set $\mathcal{S}(Q)$ of size $K$, we evaluate retrieval quality using Recall@$K$ and Acc@$K$, defined as
$\text{Recall@}K=\frac{|\mathcal{V}^{*}\cap\mathcal{S}(Q)|}{|\mathcal{V}^{*}|},\qquad\text{Acc@}K=\mathbb{I}[\mathcal{V}^{*}\subseteq\mathcal{S}(Q)]$
where $\mathcal{V}^{*}$ denotes the set of ground-truth patched functions.
Recall@$K$ measures the fraction of ground-truth functions recovered within the top-$K$ retrieved results, while Acc@$K$ is a stricter metric that equals 1 only when all ground-truth functions are included in the top-$K$ set. These metrics are particularly well-suited to the multi-function setting, as they directly capture coverage under a fixed retrieval budget. For completeness, we additionally report Mean Reciprocal Rank (MRR) in App. section˜A.10, defined as the reciprocal of the rank at which any ground-truth function first appears in the retrieved list.
Hyperparameters  We analyze the effects of hyperparameters $C$, $d$, NN, and $K$ in section˜5.3. Following prior work Reddy et al. (2025); Fehr et al. (2025), we fix the retrieval budget to $K=20$ across all datasets to ensure fair comparison with baselines that are sensitive to this choice, such as embedding-based retrieval followed by LLM reranking. We select the remaining hyperparameters of SpIDER via grid search on the Python split of SWE-PolyBench, resulting in $C=5$, $d=4$, and $N=500$. These values are held constant for all experiments to avoid overfitting. In our ablation studies, we further examine how varying these parameters impacts retrieval performance as well as practical considerations such as LLM context length requirements and invocation budget. For all experiments, we use Claude Sonnet 4 with a temperature of $0.1$ to ensure consistent and reproducible evaluation.
Baselines  We compare SpIDER against representative state-of-the-art retrieval approaches spanning both dense and sparse methods. For dense retrieval, we consider SweRankEmbed-Small ZS and CodeSAGE-Small ZS, which are widely used embedding-based models for software issue localization and code retrieval Reddy et al. (2025); Zhang et al. (2024). As a sparse retrieval baseline, we include BM25 Robertson et al. (1994).
Since CodeSAGE-Small ZS is not trained for GitHub-style problem-solving tasks, we additionally evaluate a parameter-efficiently fine-tuned variant using LoRA (Hu et al., 2022) on $2000$ instances from three repositories in the SWEBench training set, denoted as CodeSAGE-Small-PEFT.
Some recent methods could not be fully evaluated. Fehr et al. (2025) did not release their embedding model or code, preventing direct comparison. LocAgent (Chen et al., 2025) is based on a time consuming and iterative LLM-driven exploration and is therefore evaluated only on SWEBench-Verified and the Python repositories of SWE-PolyBench in tables 8 and 9.
@@ 5.2 Results
Table 2: Main result. Across all embedding models and programming languages, SpIDER consistently improves Recall and Accuracy over dense retrieval alone, with gains preserved and often amplified after LLM reranking. Here, $K=20$ with $N=500$, $C=5$, and $d=4$; best results per embedding model are highlighted in light blue. SpISR denotes the spatially informed sparse retriever analogous to SpIDER. KDE plots are provided in Figs. 4 and 5 while confidence intervals are provided in table˜10.
Language Model Retrieval SWE-PolyBench SWEBench-Verified
Dense Retrieval Only Dense + LLM reranker Dense Retrieval Only
Recall@20 Accuracy@20 Recall@3 Accuracy@3 Recall@20 Accuracy@20
Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.44 0.36 0.61 0.56
DER 0.42 0.35 0.38 0.31 0.54 0.49
CodeSAGE-Small ZS SpIDER 0.27 0.21 0.23 0.18 0.30 0.27
DER 0.21 0.17 0.19 0.16 0.21 0.19
CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.32 0.25 0.55 0.49
DER 0.29 0.20 0.26 0.20 0.43 0.38
BM25 SpISR 0.33 0.26 0.30 0.23 0.45 0.41
SR 0.31 0.24 0.27 0.22 0.38 0.34